OpenStudy (darkworld131):

What is the length of the base of a triangle with an altitude of 7cm and an area of 49cm^2 A:42cm B:7cm C:21cm D:14cm none of the above

jimthompson5910 (jim_thompson5910):

Area = (base*height)/2 A = (bh)/2 49 = (b*7)/2 ... note: altitude = height = 7 cm Now solve that last equation for b to find the base

OpenStudy (anonymous):

A of triangle =base x height. Since area and height is given, so just substitute in the equation to get length of the base. A=bxh 49=b x 7 49/7 (making b the subject of the formula)=7cm, which is the length of the base of the triangle.

OpenStudy (anonymous):

its A =bxh 49=bx7 Making b the subject of the formula, so it will be 49/7 which will be 7cm, as the length of the base of the triangle

OpenStudy (darkworld131):

7 wasnt the right answer, 14 was i guess you take 49/7 and then multiply by 2and you get 14

jimthompson5910 (jim_thompson5910):

Sancy used the wrong formula to get the wrong answer of 7. Use the formula I posted above.

OpenStudy (darkworld131):

@jim_thompson5910 i did go back and redo the problem with your formula and it worked thanks

jimthompson5910 (jim_thompson5910):

that's great, yw